### v1

@@ 4 Convergence Analysis of The P-BCD Algorithm
$\alpha_{\mathcal{G}}([\mathcal{G}^{k-1}]-[\mathcal{G}^{k}])+([\mathcal{O}^{k}]-[\tilde{\mathcal{O}}^{k}])\in[\mathcal{G}^{k}]-[\tilde{\mathcal{O}}^{k}]+\partial\psi([\mathcal{G}^{k}]),$
${0}=\delta(\mathcal{L}^{k}+\mathcal{S}^{k}-\mathcal{D})+\mathcal{L}^{k}-(\mathcal{W}_{\mathbf{R}})^{-1}\mathbf{R}^{\top}([\mathcal{Y}^{k}]),$
where $[\tilde{\mathcal{O}}^{k}]=\mathbf{R}(\mathcal{L}^{k})\times_{1}[X_{1}^{k}]^{\top}\times_{2}[X_{2}^{k}]^{\top}\times_{3}[X_{3}^{k}]^{\top}$. Since we have
$\|[\mathcal{O}^{k}]-[\tilde{\mathcal{O}}^{k}]\|_{F}$ $\leq c_{3}\|[X_{1}^{k}]\|_{F}\|[X_{2}^{k}]\|_{F}\|[X_{3}^{k}]\|_{F}\|\mathcal{L}^{k}-\mathcal{L}^{k-1}\|_{F}$
$\leq c_{3}\sqrt{n_{1}n_{2}n_{3}N^{3}}\|\mathcal{L}^{k}-\mathcal{L}^{k-1}\|_{F},$
it immediately follows from Corollary 4.3 that $\lim_{k\to\infty}\|[\mathcal{O}^{k}]-[\tilde{\mathcal{O}}^{k}]\|_{F}=0$.
According to the definition of limiting subdifferential and the fact that $\varphi$ and $\psi$ are continuous functions, we can take the limits of the relations above as $k\in\mathcal{K}$ approaches $\infty$. Note that $\mathcal{Z}^{k}\to\bar{\mathcal{Z}}$, $[\tilde{\mathcal{O}}^{k}]\to[\bar{\mathcal{O}}]$ and $[\mathcal{Y}^{k}]\to[\bar{\mathcal{Y}}]$, as $k\in\mathcal{K}$ approaches $\infty$. Together using Theorem 4.1 (iii) and Corollary 4.3, we have (15a), (15e) and (15f) hold.
@@
$\Phi(\mathcal{Z}^{k})-\Phi(\mathcal{Z}^{k+1})\geq c_{4}\sum_{i=1}^{3}\left\|([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top})[H_{i}^{k}]\right\|_{F}^{2}$
@@
$\sum^{\infty}_{k=0}\sum_{i=1}^{3}\left\|([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top})[H_{i}^{k}]\right\|_{F}^{2}\leq\frac{1}{c_{4}}\Phi(\mathcal{Z}^{0}).$
@@
Since $\mathcal{Z}^{k}\to\bar{\mathcal{Z}}$ as $k\in\mathcal{K}$ approaches $\infty$, we have $[{P}_{i}^{k}]\to[\bar{P}_{i}]$, $[Q_{1}^{k}]\to[\bar{Q}_{1}]$, as $k\in\mathcal{K}$ approaches $\infty$. Also, we have
$\|[Q_{2}^{k}]-[\bar{Q}_{2}]\|_{F}$ $\leq\|[Q_{2}^{k}]-[\tilde{Q}_{2}^{k}]\|_{F}+\|[\tilde{Q}_{2}^{k}]-[\bar{Q}_{2}]\|_{F}$
$\leq c_{2}\|[X_{1}^{k+1}]-[X_{1}^{k}]\|_{F}+\|[\tilde{Q}_{2}^{k}]-[\bar{Q}_{2}]\|_{F}$
@@
$\|[Q_{3}^{k}]-[\bar{Q}_{3}]\|_{F}$ $\leq\|[Q_{3}^{k}]-[\tilde{Q}_{3}^{k}]\|_{F}+\|[\tilde{Q}_{3}^{k}]-[\bar{Q}_{3}]\|_{F}$
$\leq c_{2}\|[X_{1}^{k+1}]-[X_{1}^{k}]\|_{F}+c_{2}\|[X_{2}^{k+1}]-[X_{2}^{k}]\|_{F}+\|[\tilde{Q}_{3}^{k}]-[\bar{Q}_{3}]\|_{F},$
where $[\tilde{Q}_{2}^{k}]=\big{(}[\mathcal{G}^{k}]\times_{1}[X_{1}^{k}]\times_{3}[X^{k}_{3}]\big{)}_{[(2)]}$ and $[\tilde{Q}_{3}^{k}]=\left([\mathcal{G}^{k}]\times_{1}[X^{k}_{1}]\times_{2}[X^{k}_{2}]\right)_{[(3)]}$. Since $[\tilde{Q}_{i}^{k}]\to[\bar{Q}_{i}]$, $i=2,3$, by Theorem 4.1 (iii), we have $[Q_{i}^{k}]\to[\bar{Q}_{i}]$, $i=2,3$, as $k\in\mathcal{K}$ approaches $\infty$. Hence, $[H_{i}^{k}]\to[\bar{H}_{i}]$, $i=1,2,3$, as $k\in\mathcal{K}$ approaches $\infty$. By (22) and (23), we have (15b) and (15c) hold. Since Stiefel manifold is a compact set, we also have (15d) holds.
@@ 5 Application to HSI Denoising and Destriping
In this section, we present an application of the proposed model in (1) to HSI denoising and destriping and utilize the proposed P-BCD method given in Algorithm 1 for solving the model. In the following, we first introduce the sparsity-enhanced nonlocal low-rank tensor regularization for removing Gaussian noise, that is, the choice of $\mathbf{R}$ and $\psi$; and then we introduce the tensor $\ell_{2,p}$ norm for removing sparse noises with linear structures, that is, the choice of $\varphi$.
5.1 Sparsity-enhanced nonlocal low-rank tensor regularization
According to the spectral correlation and the spatial nonlocal self-similarity of HSIs, a clean HSI can be approximated by nonlocal low-rank tensors [24, 22]. To denoise the HSI via the nonlocal low-rank tensor regularization, the first step is to extract nonlocal similar tensors that may have low-rank features and the second step is to characterize the low-rankness of the tensor naturally.
@@ 5.1 Sparsity-enhanced nonlocal low-rank tensor regularization
First, we apply block matching to find similar blocks and then stack them into a fourth order nonlocal similar tensor. Given an HSI $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, we divide it into a total number of $N$ overlapping FBBs of size $r\times r\times I_{3}$. For the $j$-th FBB, we search within a local window for a total of $m_{2}$ FBBs that are mostly similar to the reference block based on Euclidean distance. Then the $j$-th nonlocal similar sub-tensor of order 3 of $\mathcal{L}$, denoted as $\mathbf{R}_{j}(\mathcal{L})$, can be formed by unfolding all the nonlocal similar FBBs in the $j$-th group and then stacking them together. An illustrative figure on nonlocal low-rank tensor extraction is shown in Fig. 2. As the nonlocal similar FBB sub-tensors are independent of each other, we can further stack them together into a fourth order nonlocal similar group tensor, denoted as $\mathbf{R}(\mathcal{L})$.
@@
To be precise, we present the formulations for the $j$-th nonlocal similar sub-tensor extraction operator $\mathbf{R}_{j}$ and then for the nonlocal similar tensor extraction operator $\mathbf{R}$. We define $U_{i}^{(l)}\in\mathbb{R}^{r\times I_{i}}$ be a binary matrix such that ×$\mathcal{L}\times_{1}U_{1}^{(l)}\times_{2}U_{2}^{(l)}$ is exactly the $l$-th FBB of $\mathcal{L}$. And we define $B_{l}$ as the Casorati matrix (a matrix whose columns are vectorized bands of the HSI) of the $l$-th FBB as follows
$B_{l}:=\operatorname{reshape}(\mathcal{L}\times_{1}U_{1}^{(l)}\times_{2}U_{2}^{(l)},m_{1},m_{3}),$
where $m_{1}=r^{2}$ and $m_{3}=I_{3}$. Then the extraction operator of the $j$-th nonlocal similar sub-tensor $\mathbf{R}_{j}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ can be defined by
$\mathbf{R}_{j}(\mathcal{L}):=\operatorname{reshape}([B_{l_{j1}}^{\top},B_{l_{j2}}^{\top},\dots,B_{l_{jm_{2}}}^{\top}]^{\top},m_{1},m_{2},m_{3}),$
where the indices $l_{j1},l_{j1},\dots,l_{jm_{2}}$ refer to the indices of FBBs that belongs to the $j$-th nonlocal similar group and $j=1,2,\dots,N$. Then the extraction operator of the nonlocal similar tensor $\mathbf{R}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$ is a linear map such that $[\mathbf{R}(\mathcal{L})]^{(j)}=\mathbf{R}_{j}(\mathcal{L})$.
Since the Frobenius inner product is invariant to reshaping, we have that for $\mathcal{Y}\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$,
$\langle\mathcal{Y},\mathbf{R}_{j}(\mathcal{L})\rangle_{F}=\sum_{i=1}^{m_{2}}\langle{Y}_{:i:},B_{l_{ji}}\rangle_{F}=\langle\mathbf{R}^{\top}_{j}(\mathcal{Y}),\mathcal{L}\rangle_{F},$
where $\mathbf{R}^{\top}_{j}:\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is defined by
(24) $\mathbf{R}^{\top}_{j}(\mathcal{Y}):=\sum_{i=1}^{m_{2}}\operatorname{reshape}(Y_{:i:},r,r,I_{2})\times_{1}(U_{1}^{(l_{ji})})^{\top}\times_{2}(U_{2}^{(l_{ji})})^{\top}.$
And we further have that
$\|\mathbf{R}(\mathcal{L})\|^{2}_{F}=\sum_{j=1}^{N}\|\mathbf{R}_{j}(\mathcal{L})\|_{F}^{2}=\sum_{j=1}^{N}\langle\mathcal{L},\mathbf{R}_{j}^{\top}\mathbf{R}_{j}(\mathcal{L})\rangle_{F}=\|\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\mathcal{L}\|_{F}^{2},$
where $\odot$ represents the pointwise multiplication, and each entry of $\mathcal{W}_{\mathbf{R}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ represents the number of nonlocal similar groups to which the corresponding pixel belongs. Since we assume that each pixel belongs to at least one nonlocal similar group, we have $\mathcal{W}_{\mathbf{R}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}_{++}$.
Second, we impose the tensor low-rankness on the nonlocal similar tensor. In the nonlocal sub-tensor $\mathbf{R}_{j}(\mathcal{L})$ that we construct, the first dimension indicates the spatial information, the second dimension reveals the nonlocal self-similarity, and the third dimension reflects the spectral correlation. We adopt the independent 3-D HOSVD to obtain a low-rank approximation of $\mathbf{R}(\mathcal{L})$, that is,

### v2

@@ 4 Convergence Analysis of The P-BCD Algorithm
$\alpha_{\mathcal{G}}([\mathcal{G}^{\tilde{k}-1}]-[\mathcal{G}^{\tilde{k}}])+\delta([\mathcal{O}^{\tilde{k}}]-[\tilde{\mathcal{O}}^{\tilde{k}}])\in\delta([\mathcal{G}^{\tilde{k}}]-[\tilde{\mathcal{O}}^{\tilde{k}}])+\partial\psi([\mathcal{G}^{\tilde{k}}]),$
${0}=\mathcal{L}^{\tilde{k}}+\mathcal{S}^{\tilde{k}}-\mathcal{D}+\delta\mathcal{W}_{\mathbf{R}}\odot\mathcal{L}^{\tilde{k}}-\delta\mathbf{R}^{\top}([\mathcal{Y}^{\tilde{k}}]),$
where $[\tilde{\mathcal{O}}^{\tilde{k}}]=\mathbf{R}(\mathcal{L}^{\tilde{k}})\times_{1}[X_{1}^{\tilde{k}}]^{\top}\times_{2}[X_{2}^{\tilde{k}}]^{\top}\times_{3}[X_{3}^{\tilde{k}}]^{\top}$. Since we have
$\|[\mathcal{O}^{\tilde{k}}]-[\tilde{\mathcal{O}}^{\tilde{k}}]\|_{F}$ $\leq\sqrt{c_{2}}\|[X_{1}^{\tilde{k}}]\|_{F}\|[X_{2}^{\tilde{k}}]\|_{F}\|[X_{3}^{\tilde{k}}]\|_{F}\|\mathcal{L}^{\tilde{k}}-\mathcal{L}^{\tilde{k}-1}\|_{F}$
$\leq\sqrt{c_{2}n_{1}n_{2}n_{3}N^{3}}\|\mathcal{L}^{\tilde{k}}-\mathcal{L}^{\tilde{k}-1}\|_{F},$
it immediately follows from Corollary 4.3 that $\lim_{\tilde{k}\to\infty}\|[\mathcal{O}^{\tilde{k}}]-[\tilde{\mathcal{O}}^{\tilde{k}}]\|_{F}=0$.
According to the definition of limiting subdifferential and the fact that $\varphi$ and $\psi$ are continuous functions, we can take the limits of the relations above as $\tilde{k}\in\mathcal{K}$ approaches $\infty$. Note that $\mathcal{Z}^{\tilde{k}}\to\bar{\mathcal{Z}}$, $[\tilde{\mathcal{O}}^{\tilde{k}}]\to[\bar{\mathcal{O}}]$ and $[\mathcal{Y}^{\tilde{k}}]\to[\bar{\mathcal{Y}}]$, as $\tilde{k}\in\mathcal{K}$ approaches $\infty$. Together using Theorem 4.1 (iii) and Corollary 4.3, we have (15a), (15e) and (15f) hold.
@@
$\Phi(\mathcal{Z}^{k})-\Phi(\mathcal{Z}^{k+1})\geq c_{3}\sum_{i=1}^{3}\left\|([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top})[H_{i}^{k}]\right\|_{F}^{2}$
@@
$\sum^{\infty}_{k=0}\sum_{i=1}^{3}\left\|([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top})[H_{i}^{k}]\right\|_{F}^{2}\leq\frac{1}{c_{3}}\Phi(\mathcal{Z}^{0}).$
@@
Since $\mathcal{Z}^{\tilde{k}}\to\bar{\mathcal{Z}}$ as $\tilde{k}\in\mathcal{K}$ approaches $\infty$, we have $[{P}_{i}^{\tilde{k}}]\to[\bar{P}_{i}]$, $[Q_{1}^{\tilde{k}}]\to[\bar{Q}_{1}]$, as $\tilde{k}\in\mathcal{K}$ approaches $\infty$. Also, we have
$\|[Q_{2}^{\tilde{k}}]-[\bar{Q}_{2}]\|_{F}$ $\leq\|[Q_{2}^{\tilde{k}}]-[\tilde{Q}_{2}^{\tilde{k}}]\|_{F}+\|[\tilde{Q}_{2}^{\tilde{k}}]-[\bar{Q}_{2}]\|_{F}$
$\leq c_{1}\|[X_{1}^{\tilde{k}+1}]-[X_{1}^{\tilde{k}}]\|_{F}+\|[\tilde{Q}_{2}^{\tilde{k}}]-[\bar{Q}_{2}]\|_{F}$
@@
$\|[Q_{3}^{\tilde{k}}]-[\bar{Q}_{3}]\|_{F}$ $\leq\|[Q_{3}^{\tilde{k}}]-[\tilde{Q}_{3}^{\tilde{k}}]\|_{F}+\|[\tilde{Q}_{3}^{\tilde{k}}]-[\bar{Q}_{3}]\|_{F}$
$\leq c_{1}\|[X_{1}^{\tilde{k}+1}]-[X_{1}^{\tilde{k}}]\|_{F}+c_{1}\|[X_{2}^{\tilde{k}+1}]-[X_{2}^{\tilde{k}}]\|_{F}+\|[\tilde{Q}_{3}^{\tilde{k}}]-[\bar{Q}_{3}]\|_{F},$
where $[\tilde{Q}_{2}^{\tilde{k}}]=\big{(}[\mathcal{G}^{\tilde{k}}]\times_{1}[X_{1}^{\tilde{k}}]\times_{3}[X^{\tilde{k}}_{3}]\big{)}_{[(2)]}$ and $[\tilde{Q}_{3}^{\tilde{k}}]=\left([\mathcal{G}^{\tilde{k}}]\times_{1}[X^{\tilde{k}}_{1}]\times_{2}[X^{\tilde{k}}_{2}]\right)_{[(3)]}$. Since $[\tilde{Q}_{i}^{\tilde{k}}]\to[\bar{Q}_{i}]$, $i=2,3$, by Theorem 4.1 (iii), we have $[Q_{i}^{\tilde{k}}]\to[\bar{Q}_{i}]$, $i=2,3$, as $\tilde{k}\in\mathcal{K}$ approaches $\infty$. Hence, $[H_{i}^{\tilde{k}}]\to[\bar{H}_{i}]$, $i=1,2,3$, as ${k}\in\mathcal{K}$ approaches $\infty$. By (22) and (23), we have (15b) and (15c) hold. Since Stiefel manifold is a compact set, we also have (15d) holds.
@@ 5 Application to HSI Denoising and Destriping
In this section, we present an application of the proposed model in (1) to HSI denoising and destriping and utilize the proposed P-BCD method given in Algorithm 1 for solving the model.
@@
To denoise the HSI via the low-rank tensor regularization, we extract global, local, and nonlocal tensors that may have low-rank features via the block extraction operator $\mathbf{R}$ and impose the low-rankness on the tensors. In the following, we first introduce the block extraction for global, local, and nonlocal tensors, i.e., the choices of $\mathbf{R}$, and the sparsity enhancement on low-rank regularization for removing Gaussian noise, i.e., the choice of $\psi$. Then we introduce the tensor $\ell_{2,p}$ norm for removing sparse noises with linear structures, i.e., the choice of $\varphi$. Lastly, we propose the MLTL2p method using a multi-scale low-rank tensor regularization with a two-phase scheme to increase the effectiveness and efficiency of model (1) and Algorithm 1.
5.1 Block extraction
According to the spectral correlation and spatial nonlocal self-similarity of HSIs, a clean HSI can be approximated using global, local, and nonlocal low-rank tensors [26, 28]. These tensors differ in their extraction operators $\mathbf{R}$, which we later denote as $\mathbf{R}_{g}$ for global, $\mathbf{R}_{l}$ for local, and $\mathbf{R}_{nl}$ for nonlocal.
We begin by defining several types of block extraction operators and their transpose operators. Given an HSI $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, we can divide it into a total number of $N$ overlapping or non-overlapping blocks of size $r_{1}\times r_{2}\times r_{3}$, $r_{i}\leq I_{i}$. We define the operator that extracts the $l$-th block as $\mathbf{B}_{l}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$
$\mathbf{B}_{l}(\mathcal{L}):=\mathcal{L}\times_{1}(U_{1}^{(l)})^{\top}\times_{2}(U_{2}^{(l)})^{\top}\times_{3}(U_{3}^{(l)})^{\top},$
where $U_{i}^{(l)}\in\mathbb{R}^{I_{i}\times r_{i}}$ is a binary matrix such that $\mathbf{B}_{l}(\mathcal{L})$ is exactly the $l$-th local block of $\mathcal{L}$, and $(U_{i}^{(l)})^{\top}U_{i}^{(l)}=\mathbf{I}_{r_{i}}$. The transpose of $\mathbf{B}_{l}$ can be defined by $\mathbf{B}_{l}^{\top}:\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$
$\mathbf{B}_{l}^{\top}(\mathcal{Y}):=\mathcal{Y}\times_{1}U_{1}^{(l)}\times_{2}U_{2}^{(l)}\times_{3}U_{3}^{(l)},$
which satisfies $\langle\mathcal{Y},\mathbf{B}_{l}(\mathcal{L})\rangle_{F}=\langle\mathbf{B}^{\top}_{l}(\mathcal{Y}),\mathcal{L}\rangle_{F}$ for any $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\mathcal{Y}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$. For those $N$ independent blocks, we can stack them together and form a new operator $\mathbf{B}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{r_{1}\times r_{2}\times r_{3}\times N}$ such that $[\mathbf{B}(\mathcal{L})]^{(l)}=\mathbf{B}_{l}(\mathcal{L})$.
Depending on the size of the blocks, there are several types of block extraction as follows using the formulation of $\mathbf{B}$:
  • Global extraction $\mathbf{R}_{g}$: $\mathbf{B}$ with $r_{i}=I_{i}$, $i=1,2,3$;
  • Local extraction $\mathbf{R}_{l}$: $\mathbf{B}$ with $r_{i}<I_{i}$, $i=1,2,3$;
  • Full band extraction $\mathbf{R}_{f}$: $\mathbf{B}$ with $r_{i}<I_{i}$, $i=1,2$ and $r_{3}=I_{3}$.
The tensors formed by the above block extraction operators are based on the spectral correlation of HSIs.
Next, we formulate another type of block extraction operators based on spatial nonlocal self-similarity of HSIs. After dividing the HSI into $N$ blocks using $\mathbf{B}_{l}$, we apply block matching to find similar blocks and then stack them into a fourth order nonlocal similar tensor.
@@ 5.1 Block extraction
Taking FBBs as an example, for the $j$-th FBB, we search within a local window for a total of $m_{2}$ FBBs that are mostly similar to the reference FBB based on Euclidean distance. Then the $j$-th nonlocal similar sub-tensor of order 3 of $\mathcal{L}$, denoted as $\mathbf{R}_{j}(\mathcal{L})$, can be formed by unfolding all the nonlocal similar blocks in the $j$-th group and then stacking them together. We define the Casorati matrix (a matrix whose columns are vectorized bands of the HSI) of the $l$-th FBB as follows
$\tilde{B}_{l}:=\operatorname{reshape}(\mathbf{B}_{l}(\mathcal{L}),m_{1},m_{3}),$
where $m_{1}=r_{1}r_{2}$ and $m_{3}=I_{3}$. Then the extraction operator of the $j$-th nonlocal similar sub-tensor $\mathbf{R}_{j}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}$ can be defined by
$\mathbf{R}_{j}(\mathcal{L}):=\operatorname{reshape}([\tilde{{B}}_{l_{j1}}^{\top},\tilde{{B}}_{l_{j2}}^{\top},\dots,\tilde{{B}}_{l_{jm_{2}}}^{\top}]^{\top},m_{1},m_{2},m_{3}),$
where the indices $l_{j1},l_{j1},\dots,l_{jm_{2}}$ refer to the indices of FBBs that belongs to the $j$-th nonlocal similar group and $j=1,2,\dots,N$. An illustrative figure on nonlocal low-rank tensor extraction is shown in Fig. 4. Then the transpose $\mathbf{R}^{\top}_{j}:\mathbb{R}^{m_{1}\times m_{2}\times m_{3}}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is defined by
(24) $\mathbf{R}^{\top}_{j}(\mathcal{Y}):=\sum_{i=1}^{m_{2}}\mathbf{B}_{l_{ji}}^{\top}\left(\operatorname{reshape}(Y_{:i:},r_{1},r_{2},I_{3})\right),$
which satisfies $\langle\mathcal{Y},\mathbf{R}_{j}(\mathcal{L})\rangle_{F}=\sum_{i=1}^{m_{2}}\langle{Y}_{:i:},\mathbf{B}_{l_{ji}}(\mathcal{L})\rangle_{F}=\langle\mathbf{R}^{\top}_{j}(\mathcal{Y}),\mathcal{L}\rangle_{F}$. Stacking $\mathbf{R}_{j}(\mathcal{L})$ into a fourth order nonlocal similar group tensor, the extraction operator $\mathbf{R}_{nl}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$ is defined as a linear map such that $[\mathbf{R}_{nl}(\mathcal{L})]^{(j)}=\mathbf{R}_{j}(\mathcal{L})$.
@@
Note that the $\mathbf{R}$ formulated above all satisfy the following equality
$\mathcal{W}_{\mathbf{R}}\odot\operatorname{Id}=\mathbf{R}^{\top}\mathbf{R},$
where $\odot$ represents the pointwise multiplication, and each entry of $\mathcal{W}_{\mathbf{R}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ represents the number of blocks to which the corresponding pixel belongs. For $\mathbf{R}_{g}$, $\mathbf{R}_{l}$ and $\mathbf{R}_{f}$, the weight tensor can be defined by
$\mathcal{W}_{\mathbf{R}}\odot\operatorname{Id}=\sum_{l=1}^{N}\mathbf{B}_{l}^{\top}\mathbf{B}_{l},$
and for $\mathbf{R}_{nl}$, the weight tensor can be defined by
$\mathcal{W}_{\mathbf{R}}\odot\operatorname{Id}=\sum_{j=1}^{N}\mathbf{R}_{j}^{\top}\mathbf{R}_{j}.$
Since we assume that each pixel belongs to at least one low-rank group, we have $\mathcal{W}_{\mathbf{R}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}_{++}$.
5.2 Sparsity enhanced low-rank tensor regularization
Taking $\mathbf{R}_{nl}$ as an example, in the $\mathbf{R}_{nl}(\mathcal{L})$ that we construct, the first dimension indicates the spatial information, the second dimension reveals the nonlocal self-similarity, the third dimension reflects the spectral correlation, and the fourth dimension shows no correlation. Hence, we adopt the independent 3-D HOSVD to obtain a low-rank approximation of $\mathbf{R}(\mathcal{L})$, that is,